# MOCHA: Multi-modal Objects-aware Cross-arcHitecture Alignment

### v3

@@ Appendix A.2 Dimensionality Reduction
Fig. A.12 shows the variance profile of teacher embeddings, with a fitted decay curve of the form $\sigma\simeq\frac{a}{(x+1)^{b}}+c$, confirming the typical power-law structure that motivates dimensionality reduction. The standard deviation of each channel activation in the distillation dataset is estimated and used to rescale the channel $c$ as: $\sigma_{c}\simeq\frac{18}{(c+1)^{0.47}}-0.26$ (Eq. 3 in the main document). Tables A.6, A.7 and A.5 report extensive results across all datasets and protocols for different PCA settings under three configurations:
@@
  • PCA on $\hat{u}^{\prime}_{i}$ in Personal Dataset (Tab. A.5), where PCA is applied on teacher embeddings of each personal dataset (ideal case);
  • PCA on $\hat{u}^{\prime}_{i}$ in Distillation Dataset (Tab. A.6), where PCA is applied on teacher embeddings of distillation dataset;
@@
  • PCA on $u^{\prime}_{i}$ in Distillation Dataset (Tab. A.7), where the channel normalization of Fig. A.12 is applied prior to PCA.
@@
Across all experiments, we observe that moderate compression (e.g., retaining 128128256256 dimensions) achieves competitive performance, with minimal loss relative to full-dimensional embeddings, while substantially reducing memory and compute costs. As expected, aggressive compression (e.g., 16163232 dimensions) leads to consistent drops across datasets, particularly on POD and iCubWorld. This analysis confirms that PCA acts as an effective mechanism to control embedding size while preserving sufficient discriminative information for distillation. The results validate the robustness of our approach to dimensionality reduction, supporting the choice of PCA dimension $d_{t}=512$.
Appendix A.3 Qualitative Results
Figures A.9 and A.11 present qualitative comparisons across three models—YOLO (baseline), AuXFT, and MOCHA (AuXFT)—in both the 1-shot and 5-shot personalization settings. Results are shown on three representative datasets: POD, CoRE50, and iCubWorld. Ground truth annotations are depicted in green, while the prediction with the highest confidence is shown in red. In the 1-shot setting (Figure A.9), MOCHA already improves over both YOLO and AuXFT, correctly localizing and classifying objects that other models miss or mislabel (e.g., remote2 in POD, case0 in iCubWorld). With 5-shot supervision (Figure A.11), MOCHA further enhances precision, handling subtle distinctions (e.g., pen5 vs. mug5 in CoRE50), and resolving multiple similar instances in cluttered scenes. These examples qualitatively validate MOCHA’s effectiveness in aligning student features to the teacher’s multimodal embedding space, enabling more reliable object personalization with minimal data.
Appendix A.4 Implementation Details
All experiments were run on a RHEL8 (RedHat) Unix-based machine with kernel version 6.12.9-1, equipped with 8 NVIDIA L40S GPUs (48GB of VRAM, CUDA 12.8, Driver 570.86.15), 2 AMD EPYC 9224 24-Core processors, 1.5TB of RAM, and Python version 3.11.9. Each pretraining experiment uses 4 GPUs, 24 CPU cores, and 64GB of RAM, while for few-shot training and inference, a single GPU and 6 CPU cores are sufficient. A complete MOCHA distillation (50 epochs) takes about 48 hours, which decreases to around 20 hours when distilling from the AuXFT checkpoints (20 epochs).
A.4.1 MOCHA Pseudocode
In this section, we report the pseudocode implementation of MOCHA’s feature distillation and few-shot learning procedures. Algorithm A.1 reports the steps necessary to distill visual-language knowledge into an efficient vision-only detector, while Algorithms A.2 and A.3 explain how to train MOCHA’s few-shot module and perform inference, respectively. Note that # marks a comment.
Algorithm A.1 Pseudocode for MOCHA’s feature distillation.
@@ A.4.1 MOCHA Pseudocode
17:     $l\leftarrow l+\mathcal{L}_{\mathrm{dist}}(f^{\prime}_{A,i},u_{i})/n$   # $n$ = num of boxes
@@
Appendix A.5 MOCHA Limitations
The improvements brought about by MOCHA are substantial, but we wish to highlight some limitations that we consider potential directions for future research. First, due to the nature of the MOCHA architecture, the final task is constrained to object detection, which may limit applicability in other domains. Second, while inference requires minimal overhead, the reliance on foundation models during training incurs significant computational cost, which is undesirable in some contexts. Finally, the prototypical network in our method is inherently inefficient at large scales, meaning that performance may degrade when a large number of personal object instances are involved.
Algorithm A.2 Pseudocode for MOCHA’s FSL training.
@@ Appendix A.5 MOCHA Limitations
Algorithm A.3 Pseudocode for MOCHA’s FSL inference.
0: Input image $X$
0: Frozen distilled model $\mathrm{m}_{S}=\mathrm{l}_{S}\circ\mathrm{g}_{S}$, frozen translator $\mathrm{t}_{S}$, trained prototype classifier $\mathrm{p}(\cdot)$
[…]
@@
YOLO
@@
Figure A.9: Qualitative results in the 1-shot setting. Ground truth in green, prediction with the highest confidence in red (class names shown refer to personal class labels in each dataset). YOLO refers to the baseline. *: refers to MOCHA (AuXFT).
YOLO
@@
Figure A.11: Qualitative results in the 5-shot setting. Ground truth in green, prediction with the highest confidence in red (class names shown refer to personal class labels in each dataset). YOLO refers to the baseline. *: refers to MOCHA (AuXFT).
Figure A.12: Fitting the PCA curve $\sigma\simeq\frac{a}{(x+1)^{b}}+c$, a=18,b=0.47,c=0.26a=18,b=0.47,c=-0.26.
Model PerSeg POD CORe50 iCubWorld Avg
1 shot 1 shot 5 shot 1 shot 5 shot 1 shot 5 shot
PCA 1024 (4.5x less channels) 94.2 $\pm$ 1.7 67.9 $\pm$ 3.4 80.6 $\pm$ 0.0 60.1 $\pm$ 4.6 80.8 $\pm$ 4.3 59.3 $\pm$ 3.1 81.2 $\pm$ 2.3 74.9
PCA 512 (9x less channels) 94.3 $\pm$ 1.7 67.9 $\pm$ 3.5 80.4 $\pm$ 0.0 60.3 $\pm$ 4.6 80.6 $\pm$ 4.3 59.3 $\pm$ 3.1 80.9 $\pm$ 2.3 74.8
PCA 384 (12x less channels) 94.4 $\pm$ 1.7 67.5 $\pm$ 3.4 80.0 $\pm$ 0.0 60.4 $\pm$ 4.6 80.6 $\pm$ 4.3 59.3 $\pm$ 3.0 80.8 $\pm$ 2.2 74.7
PCA 256 (18x less channels) 94.2 $\pm$ 1.6 67.6 $\pm$ 3.3 79.6 $\pm$ 0.0 60.6 $\pm$ 4.6 80.5 $\pm$ 4.5 59.2 $\pm$ 3.1 80.2 $\pm$ 2.2 74.6
PCA 128 (36x less channels) 92.9 $\pm$ 1.5 66.4 $\pm$ 3.1 79.4 $\pm$ 0.0 60.5 $\pm$ 4.6 79.8 $\pm$ 4.3 57.5 $\pm$ 3.1 77.6 $\pm$ 2.4 73.4
PCA 64 (72x less channels) 88.0 $\pm$ 2.9 63.9 $\pm$ 3.4 79.9 $\pm$ 0.0 59.0 $\pm$ 4.5 77.1 $\pm$ 4.6 52.7 $\pm$ 3.3 71.2 $\pm$ 2.4 70.3
PCA 32 (144x less channels) 81.5 $\pm$ 2.8 60.0 $\pm$ 3.5 70.6 $\pm$ 0.0 53.5 $\pm$ 4.7 71.4 $\pm$ 5.0 45.9 $\pm$ 3.1 62.1 $\pm$ 2.7 63.6
Table A.5: PCA on $\hat{u}^{\prime}_{i}$ in Personal Dataset (mAP/mAcc ± std). The last column (Avg) reports the average across the 7 columns. Overall best in bold.
Model PerSeg POD CORe50 iCubWorld Avg
1 shot 1 shot 5 shot 1 shot 5 shot 1 shot 5 shot
PCA 1024 (4.5x less channels) 94.4 $\pm$ 1.7 67.9 $\pm$ 3.5 80.3 $\pm$ 0.0 59.7 $\pm$ 4.5 80.3 $\pm$ 4.3 59.2 $\pm$ 3.3 81.2 $\pm$ 2.2 74.7
PCA 512 (9x less channels) 94.2 $\pm$ 1.6 67.8 $\pm$ 3.5 80.1 $\pm$ 0.0 59.8 $\pm$ 4.5 80.1 $\pm$ 4.3 58.9 $\pm$ 3.3 80.8 $\pm$ 2.2 74.5
PCA 384 (12x less channels) 93.7 $\pm$ 1.7 67.7 $\pm$ 3.4 80.4 $\pm$ 0.0 59.8 $\pm$ 4.6 79.9 $\pm$ 4.4 58.6 $\pm$ 3.3 80.3 $\pm$ 2.3 74.3
PCA 256 (18x less channels) 93.0 $\pm$ 1.6 67.4 $\pm$ 3.5 79.2 $\pm$ 0.0 59.3 $\pm$ 4.6 79.6 $\pm$ 4.4 57.6 $\pm$ 3.3 79.3 $\pm$ 2.2 73.6
PCA 128 (36x less channels) 90.4 $\pm$ 2.0 65.9 $\pm$ 3.1 78.6 $\pm$ 0.0 56.7 $\pm$ 4.7 77.2 $\pm$ 4.5 54.3 $\pm$ 3.3 75.3 $\pm$ 2.3 71.2
PCA 64 (72x less channels) 85.8 $\pm$ 3.1 62.5 $\pm$ 3.3 74.8 $\pm$ 0.0 51.9 $\pm$ 4.6 72.2 $\pm$ 4.7 48.4 $\pm$ 3.2 67.9 $\pm$ 2.5 66.2
PCA 32 (144x less channels) 79.3 $\pm$ 3.2 57.0 $\pm$ 3.6 72.1 $\pm$ 0.0 44.8 $\pm$ 4.7 63.4 $\pm$ 5.1 43.7 $\pm$ 3.2 60.7 $\pm$ 2.6 60.1
Table A.6: PCA on $\hat{u}^{\prime}_{i}$ in Distillation Dataset (mAP/mAcc ± std). The last column (Avg) reports the average across the 7 columns. Overall best in bold.
Model PerSeg POD CORe50 iCubWorld Avg
1 shot 1 shot 5 shot 1 shot 5 shot 1 shot 5 shot
PCA 1024 (4.5x less channels) 96.6 $\pm$ 1.3 68.5 $\pm$ 3.7 81.7 $\pm$ 0.0 58.9 $\pm$ 4.3 80.6 $\pm$ 3.9 61.2 $\pm$ 3.6 83.4 $\pm$ 2.2 75.8
PCA 512 (9x less channels) 96.9 $\pm$ 1.1 68.9 $\pm$ 3.7 81.9 $\pm$ 0.0 59.6 $\pm$ 4.3 80.6 $\pm$ 4.0 63.7 $\pm$ 3.4 85.5 $\pm$ 2.1 76.7
PCA 384 (12x less channels) 96.4 $\pm$ 1.2 68.5 $\pm$ 3.7 83.3 $\pm$ 0.0 60.4 $\pm$ 4.4 80.6 $\pm$ 4.2 64.0 $\pm$ 3.2 86.0 $\pm$ 2.0 77.0
PCA 256 (18x less channels) 95.2 $\pm$ 1.3 68.6 $\pm$ 3.6 83.2 $\pm$ 0.0 60.9 $\pm$ 4.4 81.0 $\pm$ 4.2 64.1 $\pm$ 3.3 86.0 $\pm$ 1.9 77.0
PCA 128 (36x less channels) 92.2 $\pm$ 1.9 67.7 $\pm$ 3.7 79.3 $\pm$ 0.0 59.9 $\pm$ 4.4 79.6 $\pm$ 4.3 62.6 $\pm$ 3.2 84.2 $\pm$ 1.9 75.1
PCA 64 (72x less channels) 87.9 $\pm$ 2.2 65.7 $\pm$ 3.6 75.7 $\pm$ 0.0 55.3 $\pm$ 4.5 74.5 $\pm$ 4.7 55.0 $\pm$ 3.1 76.4 $\pm$ 2.4 70.1
PCA 32 (144x less channels) 80.2 $\pm$ 3.2 60.6 $\pm$ 3.8 74.6 $\pm$ 0.0 46.4 $\pm$ 4.6 64.7 $\pm$ 5.5 48.0 $\pm$ 3.2 68.0 $\pm$ 2.8 63.2
Table A.7: PCA on $u^{\prime}_{i}$ in Distillation Dataset (mAP/mAcc ± std). The last column (Avg) reports the average across the 7 columns. Overall best in bold.

### v4

@@ A.3.3 Dimensionality Reduction
Fig. 8 shows the variance profile of teacher embeddings, with a fitted decay curve of the form $\sigma\simeq\frac{a}{(x+1)^{b}}+c$, confirming the typical power-law structure that motivates dimensionality reduction. The standard deviation of each channel activation in the distillation dataset is estimated and used to rescale the channel $c$ as: $\sigma_{c}\simeq\frac{18}{(c+1)^{0.47}}-0.26$ (Eq. 3 in the main document). Tables 10, 11 and 9 report extensive results across all datasets and protocols for different PCA settings under three configurations:
@@
  • PCA on $\hat{u}^{\prime}_{i}$ in Personal Dataset (Tab. 9), where PCA is applied on teacher embeddings of each personal dataset (ideal case);
  • PCA on $\hat{u}^{\prime}_{i}$ in Distillation Dataset (Tab. 10), where PCA is applied on teacher embeddings of distillation dataset;
@@
  • PCA on $u^{\prime}_{i}$ in Distillation Dataset (Tab. 11), where the channel normalization of Fig. 8 is applied prior to PCA.
@@
Across all experiments, we observe that moderate compression (e.g., retaining 128128256256 dimensions) achieves competitive performance, with minimal loss relative to full-dimensional embeddings, while substantially reducing memory and compute costs. As expected, aggressive compression (e.g., 16163232 dimensions) leads to consistent drops across datasets, particularly on POD and iCubWorld. Overall, this analysis confirms that PCA acts as an effective mechanism to control embedding size while preserving sufficient discriminative information for distillation. The results validate the robustness of our approach to dimensionality reduction, supporting the choice of PCA dimension $d_{t}=512$.
Algorithm 1 Pseudocode for MOCHA’s feature distillation.
@@
17:      $l\leftarrow l+\mathcal{L}_{\mathrm{dist}}(f^{\prime}_{A,i},u_{i})/n$ $n$ = num of boxes
@@
Algorithm 2 Pseudocode for MOCHA’s FSL training.
@@
Algorithm 3 Pseudocode for MOCHA’s FSL inference.
0:  Frozen distilled model $\mathrm{m}_{S}=\mathrm{l}_{S}\circ\mathrm{g}_{S}$, frozen translator $\mathrm{t}_{S}$, trained prototype classifier $\mathrm{p}(\cdot)$, image $X$
1:  $\{F_{j}\}\leftarrow\mathrm{g}_{S}(X)$ # Extract multiscale features
2:  $\hat{\mathcal{Y}}\leftarrow\mathrm{l}_{S}(\{F_{j}\})$ # Compute detection predictions for general classes
@@
A.4 Implementation Details
All experiments were run on a RHEL8 (RedHat) Unix-based machine with kernel version 6.12.9-1, equipped with 8 NVIDIA L40S GPUs (48GB of VRAM, CUDA 12.8, Driver 570.86.15), 2 AMD EPYC 9224 24-Core processors, 1.5TB of RAM, and Python version 3.11.9. Each pretraining experiment uses 4 GPUs, 24 CPU cores, and 64GB of RAM, while for few-shot training and inference, a single GPU and 6 CPU cores are sufficient. A complete MOCHA distillation (50 epochs) takes about 28 hours, which decreases to around 11 hours when distilling from the AuXFT checkpoints (20 epochs). Fine-grained information on all packages used in the environment is available as a requirements.txt file in the code repository.
A.4.1 MOCHA Pseudocode
Algorithm 1 reports the steps necessary to distill visual-language knowledge into an efficient vision-only detector, while Algorithms 2 and 3 explain how to train MOCHA’s few-shot module and perform inference, respectively. Note that lines beginning with the symbol # denote comments.
A.5 Qualitative Results
Figures 10 and 12 present qualitative comparisons across three models—YOLO (baseline), AuXFT, and MOCHA (AuXFT)—in both the 1-shot and 5-shot personalization settings. Results are shown on three representative datasets: POD, CoRE50, and iCubWorld. Ground truth annotations are depicted in green, while the prediction with the highest confidence is shown in red. In the 1-shot setting (Fig. 10), MOCHA already improves over both YOLO and AuXFT, correctly localizing and classifying objects that other models miss or mislabel (e.g., remote2 in POD, case0 in iCubWorld). With 5-shot supervision (Fig. 12), MOCHA further enhances precision, handling subtle distinctions (e.g., pen5 vs. mug5 in CoRE50), and resolving multiple similar instances in cluttered scenes. These examples qualitatively validate MOCHA’s effectiveness in aligning student features to the teacher’s multimodal embedding space, enabling more reliable object personalization with minimal data.
A.6 Limitations
While MOCHA brings notable advances, several limitations remain that open directions for future research. First, MOCHA is inherently tailored to object detection, which may restrict its applicability to other tasks. Second, although inference incurs minimal overhead, the dependence on large foundation models during training introduces considerable computational cost, which may be impractical in certain scenarios. Finally, the prototypical network component becomes inefficient at scale, leading to potential performance degradation when handling a large number of personalized object instances.
Figure 8: Fitting the PCA curve $\sigma\simeq\frac{a}{(x+1)^{b}}+c$, a=18,b=0.47,c=0.26a=18,b=0.47,c=-0.26.
Model PerSeg POD CORe50 iCubWorld Avg
1 shot 1 shot 5 shot 1 shot 5 shot 1 shot 5 shot
PCA @ $1024$ dims ($4.5\times$ compression) 94.2 $\pm 1.7$ 67.9 $\pm 3.4$ 80.6 $\pm 0.0$ 60.1 $\pm 4.6$ 80.8 $\pm 4.3$ 59.3 $\pm 3.1$ 81.2 $\pm 2.3$ 74.9
PCA @ $512$ dims ($9\times$ compression) 94.3 $\pm 1.7$ 67.9 $\pm 3.5$ 80.4 $\pm 0.0$ 60.3 $\pm 4.6$ 80.6 $\pm 4.3$ 59.3 $\pm 3.1$ 80.9 $\pm 2.3$ 74.8
PCA @ $384$ dims ($12\times$ compression) 94.4 $\pm 1.7$ 67.5 $\pm 3.4$ 80.0 $\pm 0.0$ 60.4 $\pm 4.6$ 80.6 $\pm 4.3$ 59.3 $\pm 3.0$ 80.8 $\pm 2.2$ 74.7
PCA @ $256$ dims ($18\times$ compression) 94.2 $\pm 1.6$ 67.6 $\pm 3.3$ 79.6 $\pm 0.0$ 60.6 $\pm 4.6$ 80.5 $\pm 4.5$ 59.2 $\pm 3.1$ 80.2 $\pm 2.2$ 74.6
PCA @ $128$ dims ($36\times$ compression) 92.9 $\pm 1.5$ 66.4 $\pm 3.1$ 79.4 $\pm 0.0$ 60.5 $\pm 4.6$ 79.8 $\pm 4.3$ 57.5 $\pm 3.1$ 77.6 $\pm 2.4$ 73.4
PCA @ $64$ dims ($72\times$ compression) 88.0 $\pm 2.9$ 63.9 $\pm 3.4$ 79.9 $\pm 0.0$ 59.0 $\pm 4.5$ 77.1 $\pm 4.6$ 52.7 $\pm 3.3$ 71.2 $\pm 2.4$ 70.3
PCA @ $32$ dims ($144\times$ compression) 81.5 $\pm 2.8$ 60.0 $\pm 3.5$ 70.6 $\pm 0.0$ 53.5 $\pm 4.7$ 71.4 $\pm 5.0$ 45.9 $\pm 3.1$ 62.1 $\pm 2.7$ 63.6
Table 9: PCA on $\hat{u}^{\prime}_{i}$ in Personal Dataset (mAP/mAcc ± std). Last column (Avg) reports average across the 7 columns. Best in bold.
Model PerSeg POD CORe50 iCubWorld Avg
1 shot 1 shot 5 shot 1 shot 5 shot 1 shot 5 shot
PCA @ $1024$ dims ($4.5\times$ compression) 94.4 $\pm 1.7$ 67.9 $\pm 3.5$ 80.3 $\pm 0.0$ 59.7 $\pm 4.5$ 80.3 $\pm 4.3$ 59.2 $\pm 3.3$ 81.2 $\pm 2.2$ 74.7
PCA @ $512$ dims ($9\times$ compression) 94.2 $\pm 1.6$ 67.8 $\pm 3.5$ 80.1 $\pm 0.0$ 59.8 $\pm 4.5$ 80.1 $\pm 4.3$ 58.9 $\pm 3.3$ 80.8 $\pm 2.2$ 74.5
PCA @ $384$ dims ($12\times$ compression) 93.7 $\pm 1.7$ 67.7 $\pm 3.4$ 80.4 $\pm 0.0$ 59.8 $\pm 4.6$ 79.9 $\pm 4.4$ 58.6 $\pm 3.3$ 80.3 $\pm 2.3$ 74.3
PCA @ $256$ dims ($18\times$ compression) 93.0 $\pm 1.6$ 67.4 $\pm 3.5$ 79.2 $\pm 0.0$ 59.3 $\pm 4.6$ 79.6 $\pm 4.4$ 57.6 $\pm 3.3$ 79.3 $\pm 2.2$ 73.6
PCA @ $128$ dims ($36\times$ compression) 90.4 $\pm 2.0$ 65.9 $\pm 3.1$ 78.6 $\pm 0.0$ 56.7 $\pm 4.7$ 77.2 $\pm 4.5$ 54.3 $\pm 3.3$ 75.3 $\pm 2.3$ 71.2
PCA @ $64$ dims ($72\times$ compression) 85.8 $\pm 3.1$ 62.5 $\pm 3.3$ 74.8 $\pm 0.0$ 51.9 $\pm 4.6$ 72.2 $\pm 4.7$ 48.4 $\pm 3.2$ 67.9 $\pm 2.5$ 66.2
PCA @ $32$ dims ($144\times$ compression) 79.3 $\pm 3.2$ 57.0 $\pm 3.6$ 72.1 $\pm 0.0$ 44.8 $\pm 4.7$ 63.4 $\pm 5.1$ 43.7 $\pm 3.2$ 60.7 $\pm 2.6$ 60.1
Table 10: PCA on $\hat{u}^{\prime}_{i}$ in Distillation Dataset (mAP/mAcc ± std). Last column (Avg) reports average across the 7 columns. Best in bold.
Model PerSeg POD CORe50 iCubWorld Avg
1 shot 1 shot 5 shot 1 shot 5 shot 1 shot 5 shot
PCA @ $1024$ dims ($4.5\times$ compression) 96.6 $\pm 1.3$ 68.5 $\pm 3.7$ 81.7 $\pm 0.0$ 58.9 $\pm 4.3$ 80.6 $\pm 3.9$ 61.2 $\pm 3.6$ 83.4 $\pm 2.2$ 75.8
PCA @ $512$ dims ($9\times$ compression) 96.9 $\pm 1.1$ 68.9 $\pm 3.7$ 81.9 $\pm 0.0$ 59.6 $\pm 4.3$ 80.6 $\pm 4.0$ 63.7 $\pm 3.4$ 85.5 $\pm 2.1$ 76.7
PCA @ $384$ dims ($12\times$ compression) 96.4 $\pm 1.2$ 68.5 $\pm 3.7$ 83.3 $\pm 0.0$ 60.4 $\pm 4.4$ 80.6 $\pm 4.2$ 64.0 $\pm 3.2$ 86.0 $\pm 2.0$ 77.0
PCA @ $256$ dims ($18\times$ compression) 95.2 $\pm 1.3$ 68.6 $\pm 3.6$ 83.2 $\pm 0.0$ 60.9 $\pm 4.4$ 81.0 $\pm 4.2$ 64.1 $\pm 3.3$ 86.0 $\pm 1.9$ 77.0
PCA @ $128$ dims ($36\times$ compression) 92.2 $\pm 1.9$ 67.7 $\pm 3.7$ 79.3 $\pm 0.0$ 59.9 $\pm 4.4$ 79.6 $\pm 4.3$ 62.6 $\pm 3.2$ 84.2 $\pm 1.9$ 75.1
PCA @ $64$ dims ($72\times$ compression) 87.9 $\pm 2.2$ 65.7 $\pm 3.6$ 75.7 $\pm 0.0$ 55.3 $\pm 4.5$ 74.5 $\pm 4.7$ 55.0 $\pm 3.1$ 76.4 $\pm 2.4$ 70.1
PCA @ $32$ dims ($144\times$ compression) 80.2 $\pm 3.2$ 60.6 $\pm 3.8$ 74.6 $\pm 0.0$ 46.4 $\pm 4.6$ 64.7 $\pm 5.5$ 48.0 $\pm 3.2$ 68.0 $\pm 2.8$ 63.2
Table 11: PCA on $u^{\prime}_{i}$ in Distillation Dataset (mAP/mAcc ± std). Last column (Avg) reports average across the 7 columns. Best in bold.
YOLOv8n
@@ A.6 Limitations
Figure 10: Qualitative results in the 1-shot setting. Ground truth in green, prediction with the highest confidence in red (class names shown refer to personal class labels in each dataset). YOLOv8n refers to the baseline. *: refers to MOCHA (AuXFT).
YOLOv8n
@@
Figure 12: Qualitative results in the 5-shot setting. Ground truth in green, prediction with the highest confidence in red (class names shown refer to personal class labels in each dataset). YOLOv8n refers to the baseline. *: refers to MOCHA (AuXFT).